OpenStudy (anonymous):

Which equation does the graph of the systems of equations solve? two linear functions intersecting at 2, negative 3 one halfx + 4 = 2x − 1 −one halfx + 4 = −2x − 1 −one halfx − 4 = 2x − 1 one halfx − 4 = −2x + 1

OpenStudy (anonymous):

one halfx + 4 = 2x − 1 −one halfx + 4 = −2x − 1 −one halfx − 4 = 2x − 1 one halfx − 4 = −2x + 1

OpenStudy (anonymous):

ANYONE HELP ME WITH THIS PROMBLEM PLEASE